OpenStudy (anonymous):

Help? Explain? 3x - 9y = 3 6x - 3y = -24

OpenStudy (anonymous):

These are linear equations what did you want to do with them?

OpenStudy (anonymous):

It's a system, I have to solve them but I can't figure out how to do it.

OpenStudy (anonymous):

you can solve using elimination: multiply the first equation by 2 to eliminate the x's:\[~~~~~2~(3x-9y=3)~~~~~~\implies~~~~~~~~6x-18y=6\]\[\underline{-~(6x-3y=-24)}~~~~~\implies~~-\underline{(6x-3y=-24)}\]\[~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~0-15y=30\]now solve for y:\[-15y=30\]divide by -15:\[y=-2\] knowing the value of y we could now use substitution to solve for x: pick either of the two equations and plug in the value of y\[3x-9(-2)=3\]distribute:\[3x+18=3\]subtract 18:\[3x=-15\]divide by 3:\[x=-5\] so: \(x=-5\) and \(y=-2\) hope this helps! :)

OpenStudy (anonymous):

So much! You helped me understand it better too, thank you so much!

OpenStudy (anonymous):

No problem, anytime! :)